OpenStudy (anonymous):

Lacey decides to take her five year old son to the history museum. The price for the child's ticket is 2.75 dollars less than the price for the adult's ticket. If you represent the price for the child's ticket using the variable "x", how would you write the algebraic expression for the adult's ticket price? 2.75x x2.75 x + 2.75 PLEASE HELP

OpenStudy (anonymous):

2nd option is x over 2.75

myininaya (myininaya):

x+2.75

OpenStudy (anonymous):

x+2.75

myininaya (myininaya):

A=adult C=child C=A-2.75 They want us to represent C with x x=A-2.75 we want to know how much the Adult price is solve we solve for A x+2.75=A A=x+2.75